# Vojta’s abc conjecture for entire curves in toric varieties highly ramified over the boundary

### v1

@@ 3.3. The moving GCD theorem for parabolic Riemann surfaces
Let ${\bf u}=(u_{1},\ldots,u_{n})$. Suppose that the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$. Then, by Theorem 15, we have
(3.26) $N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)$ $\leq_{\rm exc}\left(\frac{M^{\prime}}{M}+\frac{w}{u}-1+\frac{\epsilon}{M}\right)mn\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)$
(3.27) $+c^{\prime}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1).$
Let $c_{1}=4c^{\prime\prime}\epsilon^{-1}$. Then (3.21) reads
$c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1)\leq\frac{\epsilon}{3}\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r).$
Together with (3.24) and (3.25), we derive from (3.26) and the above inequality that
$N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)+c^{\prime}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
Finally, if the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly dependent over $V_{F,G}(Mb+1)$, i.e. there is a non-trivial relation
(3.28) $\sum_{\mathbf{i}}\alpha_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}=0,$
@@
where the sum runs over those $\mathbf{i}\in\mathbb{Z}_{\geq 0}^{n}$ and $|\mathbf{i}|\leq m$, and $\alpha_{\mathbf{i}}\in V_{F,G}(Mb+1)$ for each $\mathbf{i}$ such that $\alpha_{\mathbf{i}_{0}}\neq 0$ for some $\mathbf{i}_{0}$. Then we have
(3.29) $\sum_{\mathbf{i}\neq\mathbf{i}_{0}}\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}}=-1.$
Since $\alpha_{\mathbf{i}}\in V_{F,G}(Mb+1)$, we have
(3.30) $T_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}},\mathcal{Y}}(r)\leq T_{\alpha_{\mathbf{i}},\mathcal{Y}}(r)+T_{\alpha_{\mathbf{i_{0}}},\mathcal{Y}}(r)\leq 2(Mb+1)(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
By Theorem 12 and (3.30), the relation (3.29) implies
$T_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\mathcal{Y}}(r)\leq_{\operatorname{exc}}C_{2}\cdot(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
where $C_{2}=4\binom{n+m}{n}(Mb+1)$ and $c_{3}={\binom{n+m}{n}}\left({\binom{n+m}{n}}+1\right)$. Then
(3.31) $T_{\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\mathcal{Y}}(r)$ $\leq T_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\mathcal{Y}}(r)+T_{\frac{\alpha_{\mathbf{i}_{0}}}{\alpha_{\mathbf{i}}},\mathcal{Y}}(r)$
(3.32) $\leq_{\rm exc}c_{2}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
where $c_{2}=C_{2}+2(Mb+1)$. ∎
@@
For the convenience of later application, we state the following result for $n=1$, which can be easily obtained from the theory of resultant (see [19, Proposition 10]).
@@
Let $K$ be a subfield of the field of meromorphic functions on $\mathcal{Y}$. Let $F,G\in K[x_{0},x_{1}]$ be nonconstant coprime homogeneous polynomials. Then there exists a positive constant $c$ such that for any holomorphic functions $g_{0}$ and $g_{1}$ on ${\mathcal{Y}}$ with no common zeros, the following holds:
$N_{\rm gcd,\mathcal{Y}}(F(g_{0},g_{1}),G(g_{0},g_{1}),r)\leq c(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
@@ 4. The abc theorem for parabolic Riemann surfaces
We use the above key theorem (Theorem 16) in the following manner: Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are holomorphic functions on ${\mathcal{Y}}$ without zeros (i.e. $u_{i}$ are the units on ${\mathcal{Y}}$). Let $\mathbb{K}_{\mathbf{u}}$ be the subfield of the meromorphic function field such that, for any element $a\in\mathbb{K}_{\mathbf{u}}$, there exists a nonnegative real constant $c_{a}$ (which may depend on $a$) satisfying
$T_{a,{\mathcal{Y}}}(r)\leq_{\rm exc}c_{a}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\}).$
Notably, in our application, the constant $c_{a}$ can be effectively computed. We will use Theorem 16 for polynomials in the ring $\mathbb{K}_{\mathbf{u}}[x_{1},\ldots,x_{n}]$. Furthermore, we observe that $u_{i}^{\prime}/u_{i}\in\mathbb{K}_{\mathbf{u}}$ for $1\leq i\leq n$, since
(4.1) $T_{u_{i}^{\prime}/u_{i},{\mathcal{Y}}}(r)\leq_{\rm exc}{1\over 2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+2\log T_{u_{i},{\mathcal{Y}}}(r)$
by Lemma 10.
@@
(4.2) $D_{\mathbf{u}}(F)(\mathbf{x}):=\sum_{\mathbf{i}}\frac{(a_{\mathbf{i}}\mathbf{u}^{\mathbf{i}})^{\prime}}{\mathbf{u}^{\mathbf{i}}}\mathbf{x}^{\mathbf{i}}=\sum_{\mathbf{i}}\left(a_{\mathbf{i}}\cdot\sum_{j=1}^{n}i_{j}\frac{u_{j}^{\prime}}{u_{j}}\right)\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}].$
@@
Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are the units on ${\mathcal{Y}}$. Let $G$ be a non-constant polynomial in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ with no monomial factors and no repeated factors. Then $F:=D_{\mathbf{u}}(G)$ and $G$ are coprime in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ unless there exists a non-trivial tuple of integers $(m_{1},\ldots,m_{n})$ with $\sum_{i=1}^{n}|m_{i}|\leq 2\deg G$ such that $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},{\mathcal{Y}}}\leq T_{G,{\mathcal{Y}}}(r).$
@@
The first step of the proof of [9, Lemma 5.1] shows that we may assume that $G$ is irreducible. Write $G=\sum_{\mathbf{i}}b_{\mathbf{i}}\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$, which contains at least two distinct terms, since $G$ has no monomial factors. Then $D_{\mathbf{u}}(G)/G\in\mathbb{C}^{*}$ since $D_{\mathbf{u}}(G)$ is not zero and $\deg D_{\mathbf{u}}(G)=\deg G$. Comparing the coefficients of $G$ and $D_{\mathbf{u}}(G)$, for nonzero $b_{\mathbf{i}}$ and $b_{\mathbf{j}}$ in $\mathbb{K}_{\mathbf{u}}$ ($\mathbf{i}\neq\mathbf{j}$), we have from (4.2) that
$\frac{(b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}})^{\prime}}{b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}}=\frac{(b_{\mathbf{j}}\mathbf{u}^{\mathbf{j}})^{\prime}}{b_{\mathbf{j}}\mathbf{u}^{\mathbf{j}}},$
where $\mathbf{i}=(i_{1},\ldots,i_{n})$, and $\mathbf{j}=(j_{1},\ldots,j_{n})$. It implies that $\frac{b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}}{b_{\mathbf{j}}\mathbf{u}^{\mathbf{j}}}\in\mathbb{C}$. Therefore,
$T_{u_{1}^{i_{1}-j_{1}}\cdots u_{n}^{i_{n}-j_{n}},\mathcal{Y}}\leq T_{G,{\mathcal{Y}}}(r).$
@@
Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are units on ${\mathcal{Y}}$. Let $G$ be a non-constant polynomial in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ with no monomial factors and no repeated factors. Then, for any $\epsilon>0$, there exists a positive integer $m$ and a positive constants $\tilde{c}_{1}$ and $\tilde{c}_{2}$ depending only on $\epsilon$, $n$, and the degree and the coefficients of $G$, but independent of $\mathbf{u}$, such that the following holds: If
(4.5) $\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)\geq\tilde{c}_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
then we have either
(4.6) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}\tilde{c}_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq j\leq n}\{\log T_{u_{j},{\mathcal{Y}}}(r)\})$
for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum_{i=0}^{n}|m_{i}|\leq 2m$, or
$N_{G(\mathbf{u}),{\mathcal{Y}}}(0,r)-N^{(1)}_{G(\mathbf{u}),{\mathcal{Y}}}(0,r)\leq_{\rm exc}\epsilon\max_{1\leq j\leq n}\{T_{u_{j},{\mathcal{Y}}}(r)\}.$
@@
(4.7) $N_{\gcd,{\mathcal{Y}}}(G(\mathbf{u}),D_{\mathbf{u}}(G)(\mathbf{u}),r)\geq N_{G(\mathbf{u}),{\mathcal{Y}}}(0,r)-N^{(1)}_{G(\mathbf{u}),{\mathcal{Y}}}(0,r).$
Note that there exists a nonnegative real $c_{G}$ such that
(4.8) $T_{G,{\mathcal{Y}}}(r)\leq_{\rm exc}c_{G}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\}),$
since the coefficients of $G$ are in $\mathbb{K}_{\mathbf{u}}$. Furthermore,
(4.9) $T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r)$ $\leq 2T_{G,{\mathcal{Y}}}(r)+\deg G\sum_{i=1}^{n}T_{u_{i}^{\prime}/u_{i},{\mathcal{Y}}}(r)$
(4.10) $\leq_{\rm exc}(2c_{G}+n\deg G)(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\}),$
and
(4.11) $T_{G,{\mathcal{Y}}}(r)+T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r)\leq_{\rm exc}(3c_{G}+n\deg G)(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\})$
by (4.8) and (4.9). By Lemma 18 and (4.8), $G$ and $D_{\mathbf{u}}(G)$ are either coprime or
$T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}T_{G,{\mathcal{Y}}}(r)\leq_{\rm exc}c_{G}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\})$
for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum_{i=0}^{n}|m_{i}|\leq 2\deg G$, i.e., either $G$ and $D_{\mathbf{u}}(G)$ are coprime or (4.6) holds.
It remains to consider where $G$ and $D_{\mathbf{u}}(G)$ are coprime. Let $\epsilon>0$ be given, which may assume to be sufficiently small. We apply Theorem 16 and find $m$ and $c_{i}>0$, $1\leq i\leq 4$, such that if (4.5) holds with $\tilde{c}_{1}\geq c_{1}$ , then we have either
(4.12) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}c_{2}(T_{G,{\mathcal{Y}}}(r)+T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
@@
(4.13) $N_{\rm gcd,\mathcal{Y}}(G({\mathbf{u}}),D_{\mathbf{u}}(G)({\mathbf{u}}),r)\leq_{\rm exc}\frac{\epsilon}{2}\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)+c_{4}(T_{G,{\mathcal{Y}}}(r)+T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r)).$
We can conclude the proof by combining (4.11), (4.12), and (4.13), and letting
(4.14) $\tilde{c}_{1}=\max\{c_{1},2c_{4}(3c_{G}+n\deg G)\epsilon^{-1}\},~{}~{}~{}\mbox{and}~{}~{}\tilde{c}_{2}=\max\{c_{G},c_{3}+3c_{2}c_{G}+nc_{2}\deg G\}.$

### v2

@@ 3.4. The moving GCD theorem in the parabolic setting
Suppose that the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$. Then, Theorem 19 implies that, by noticing (3.44) and (3.45),
$N_{\rm gcd,\sigma}(F({\bf u}),G({\bf u}),r)$ $\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T^{M}_{u_{i},\sigma}(r)\}+{\tilde{c}}\max_{1\leq i\leq n}\{\log T_{u_{i}}(r)\}$
$~+c_{4}(T_{F,\sigma}(r)+T_{G,\sigma}(r))+c_{5}\left(\mathfrak{X}^{+}_{\sigma}(r)+\log r+c_{{1\over h},\sigma}(r)+c_{{1\over W},\sigma}(r)\right),$
where $\tilde{c}$} and $c_{i}$, $4\leq i\leq 5$, are computable positive reals.
If the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly dependent over $V_{F,G}(Mb+1)$, i.e. there is a non-trivial relation
(3.46) $\sum_{\mathbf{i}\in I}\alpha_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}=0,$
@@
where the sum runs over those $\mathbf{i}\in\mathbb{Z}_{\geq 0}^{n}$ and $|\mathbf{i}|\leq m$, $\alpha_{\mathbf{i}}\in V_{F,G}(Mb+1)$ for each $\mathbf{i}$, and $\alpha_{\mathbf{i}}\neq 0$ for $\mathbf{i}\in I$. Without loss of generality, we assume that no proper subsum of (3.46) vanishes. Then Theorem 16 implies that
$T^{M}_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\sigma}(r)$ $\leq_{\operatorname{exc}}\sum_{\mathbf{i}\in I}N_{\beta\alpha_{\mathbf{i}},\sigma}(0,r)+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
(3.47) $+c_{\frac{1}{D_{0}},\sigma}(r)+\sum_{\begin{subarray}{c}\mathbf{i}\in I\\ \mathbf{i}\neq\mathbf{i}_{0}\end{subarray}}c_{\beta\alpha_{\mathbf{i}}{\bf u}^{\bf i},\sigma}(r)+O(\max_{0\leq i\leq n}\{\log T_{u_{i}}(r)\})+O(\log(T_{F}(r)+\log F_{G}(r)),$
where $c_{3}={\binom{n+m}{n}}\left({\binom{n+m}{n}}+1\right)$. We note that
(3.48) $N_{\beta\alpha_{\mathbf{i}},\sigma}(0,r)\leq N_{\alpha_{\mathbf{i}},\sigma}(0,r)+N_{\beta,\sigma}(0,r)\leq 2(Mb+1)(T_{F,\sigma}(r)+T_{G,\sigma}(r))+c_{{1\over\alpha_{\mathbf{i}}},\sigma}(r),$
since the zero of $\beta$ comes from the pole of $\alpha_{\mathbf{i}}$ and by (3.14). Also, since $\alpha_{\mathbf{i}}\in V_{F,G}(Mb+1)$,
$T^{M}_{\frac{\alpha_{\mathbf{i}{0}}}{\alpha_{\mathbf{i}}},\sigma}(r)$ $\leq T^{M}_{1/\alpha_{\mathbf{i}},\sigma}(r)+T^{M}_{\alpha_{\mathbf{i_{0}}},\sigma}(r)\leq T^{M}_{\alpha_{\mathbf{i}},\sigma}(r)+T^{M}_{\alpha_{\mathbf{i_{0}}},\sigma}(r)+c_{{1\over\alpha_{\mathbf{i}}},\sigma}(r)$
(3.49) $\leq 2(Mb+1)(T_{F,\sigma}(r)+T_{G,\sigma}(r))+c_{{1\over\alpha_{\mathbf{i}}},\sigma}(r).$
Then, by combining (3.4), (3.48), (3.4),
$T^{M}_{\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\sigma}(r)$ $\leq T^{M}_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\sigma}(r)+T^{M}_{\frac{\alpha_{\mathbf{i}_{0}}}{\alpha_{\mathbf{i}}},\sigma}(r)\leq_{\rm exc}c_{2}(T_{F,\sigma}(r)+T_{G,\sigma}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
$+c_{\frac{1}{D_{0}},\sigma}(r)+2c_{{1\over\alpha_{\mathbf{i}}},\sigma}(r)+\sum_{\begin{subarray}{c}\mathbf{i}\in I\\ \mathbf{i}\neq\mathbf{i}_{0}\end{subarray}}c_{\beta\alpha_{\mathbf{i}}{\bf u}_{\mathbf{i}},\sigma}(r)+O(\max_{0\leq i\leq n}\{\log T_{u_{i}}(r)\}),$
where $c_{2}=4(Mb+1)$.
By Lemma 10, we can shrink $r_{j}$ small enough so that all of the quantities
$c^{+}_{1/h,\sigma}(r),\quad c^{+}_{1/W,\sigma}(r),\quad c_{1/D_{0},\sigma}(r),\quad c_{1/\alpha_{\mathbf{i}},\sigma}(r)\ \ (\mathbf{i}\in I),\quad\sum_{\begin{subarray}{c}\mathbf{i}\in I\\ \mathbf{i}\neq\mathbf{i}_{0}\end{subarray}}c_{\beta\alpha_{\mathbf{i}}\mathbf{u}^{\mathbf{i}},\,\sigma}(r)$
are bounded by $O(1)$, where the implied constant depends on $\mathbf{u}$, $F$, and $G$, but is independent of $r$. In addition, we shrink $r_{j}$ small enough such that (3.15) and the Proposition 12 for $u_{i},1\leq i\leq n$, hold, i.e. $T_{u_{i}}(r)$ and $T^{M}_{u_{i},\sigma}(r)$ differ only by $\log r+O(1)$. This proves the key theorem. ∎
@@
For the convenience of later application, we state the following result for $n=1$, which can be easily obtained from the theory of resultant (see [20, Proposition 10]).
@@
Let $K$ be a subfield of the field of meromorphic functions on $\mathbb{C}$}. Let $F,G\in K[x]$ be nonconstant coprime polynomials. Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given as above with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$. Then there exists a positive constant $c$ such that for any meromorphic function $g$ on $\mathcal{Y}$}, the following holds:
$N_{\rm gcd,\sigma}(F(g),G(g),r)\leq c(T_{F,\sigma}(r)+T_{G,\sigma}(r)).$
@@ 4. The abc theorem for parabolic Riemann surfaces
Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given as above with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$. Let ${\bf u}=(u_{1},\ldots,u_{n})$ where $u_{1},\ldots,u_{n}$ are meromorphic functions on $\mathbb{C}$} such that each restriction $u_{i}|_{\mathcal{Y}},1\leq i\leq n,$ is a unit. Let $\mathbb{K}_{\mathbf{u}}$ be the set of meromorphic function $a$ on $\mathbb{C}$} for which there exists a nonnegative constant $c_{a}$ (depending on $a$) with
(4.1) $T^{M}_{a,\sigma}(r)\leq_{\rm exc}c_{a}\left(\mathfrak{X}^{+}_{\sigma}(r)+\log r+\max_{1\leq i\leq n}\{\log T_{u_{i}}(r)\}\right),$
after shrinking radii $r_{j}$ suffciently small (depending on $a$). By (3.14) and Lemma 10, it is easy to see that $\mathbb{K}_{\mathbf{u}}$ is a subfield of the meromorphic function field. Notably, in our application, the constant $c_{a}$ can be effectively computed. We will use Theorem 20 for $K:=\mathbb{K}_{\mathbf{u}}$. Note that, by Lemma 13 and by (3.14), if $a\in\mathbb{K}_{\mathbf{u}}$, then $a^{\prime}\in\mathbb{K}_{\mathbf{u}}$, since
$T^{M}_{a^{\prime},\sigma}(r)$ $\leq T^{M}_{\frac{a^{\prime}}{a},\sigma}(r)+T^{M}_{a,\sigma}(r)\leq m_{a^{\prime}/a,\sigma}(r)+N_{a,\sigma}(0,r)+N_{a,\sigma}(\infty,r)+T^{M}_{a,\sigma}(r)$
$\leq m_{a^{\prime}/a,\sigma}(r)+T^{M}_{\frac{1}{a},\sigma}(r)+2T^{M}_{a,\sigma}(r).$
@@
(4.2) $D_{\mathbf{u}}(F)(\mathbf{x}):=\sum_{\mathbf{i}}\frac{(a_{\mathbf{i}}\mathbf{u}^{\mathbf{i}})^{\prime}}{\mathbf{u}^{\mathbf{i}}}\mathbf{x}^{\mathbf{i}}=\sum_{\mathbf{i}}\left(a_{\mathbf{i}}^{\prime}+a_{\mathbf{i}}\cdot\sum_{j=1}^{n}i_{j}\frac{u_{j}^{\prime}}{u_{j}}\right)\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}].$
@@
By Lemma 14, we see that $u_{i}^{\prime}/u_{i}\in\mathbb{K}_{\mathbf{u}}$. In the following lemma and in Theorem 23, we will identify certain monomial products of bounded degree in the $u_{i}$ that satisfy (4.1), and hence are contained in $\mathbb{K}_{\mathbf{u}}$.
@@
Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are the units on $\mathcal{Y}$}. Let $G$ be a non-constant polynomial in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ with no monomial factors and no repeated factors, and assume that at least one of its coefficients is equal to 1. Then $F:=D_{\mathbf{u}}(G)$ and $G$ are coprime in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ unless there exists a non-trivial tuple of integers $(m_{1},\ldots,m_{n})$ with $\sum_{i=1}^{n}|m_{i}|\leq 2\deg G$ such that $T^{M}_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\sigma}\leq T_{G,\sigma}(r).$
@@
Relation (4.4) implies that we may assume, as in the first step of the proof of [9, Lemma 5.1], that $G$ is irreducible. Write $G=\sum_{\mathbf{i}}b_{\mathbf{i}}\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$, which contains at least two distinct terms, since $G$ has no monomial factors. Then $D_{\mathbf{u}}(G)/G\in\mathbb{C}^{*}$ since $D_{\mathbf{u}}(G)$ is not zero and $\deg D_{\mathbf{u}}(G)=\deg G$. Comparing the coefficients of $G$ and $D_{\mathbf{u}}(G)$, for nonzero $b_{\mathbf{i}}$ and $b_{\mathbf{j}}=1$ in $\mathbb{K}_{\mathbf{u}}$ ($\mathbf{i}\neq\mathbf{j}$), we have from (4.2) that
$\frac{(b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}})^{\prime}}{b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}}=\frac{(\mathbf{u}^{\mathbf{j}})^{\prime}}{\mathbf{u}^{\mathbf{j}}},$
where $\mathbf{i}=(i_{1},\ldots,i_{n})$, and $\mathbf{j}=(j_{1},\ldots,j_{n})$. It implies that $b_{\mathbf{i}}\frac{\mathbf{u}^{\mathbf{i}}}{\mathbf{u}^{\mathbf{j}}}\in\mathbb{C}$. Therefore,
$T^{M}_{u_{1}^{j_{1}-i_{1}}\cdots u_{n}^{j_{n}-i_{n}},\sigma}\leq T_{G,\sigma}(r).$
@@
Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given as above with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$. Let ${\bf u}=(u_{1},\dots,u_{n})$, where each $u_{i}$, $1\leq i\leq n$, is a meromorphic function on $\mathbb{C}$} which is a unit on $\mathcal{Y}$}. Let $G$ be a non-constant polynomial in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ with no monomial factors and no repeated factors. Assume that at least one of the coefficients in the expansions of $G$ is equal to 1. Then, by shrinking $r_{j}$ small enough (depending on $\epsilon$, $\bf u$} and $G$), we can find a positive integer $m$ and positive constants $\tilde{c}_{1}$ and $\tilde{c}_{2}$ depending only on $\epsilon$, $n$, and on the degree and the coefficients of $G$, but independent of $\mathbf{u}$, for which the following holds: Either there exists a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum_{i=0}^{n}|m_{i}|\leq 2m$ such that
(4.5) $T^{M}_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\sigma}(r)\leq_{\rm exc}\tilde{c}_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq j\leq n}\{\log T_{u_{j}}(r)\}),$
or else
(4.6) $N_{G(\mathbf{u}),\sigma}(0,r)-N^{(1)}_{G(\mathbf{u}),\sigma}(0,r)\leq_{\rm exc}\epsilon\max_{1\leq j\leq n}\{T_{u_{j}}(r)\}+\tilde{c}_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
@@
(4.7) $N_{\gcd,\sigma}(G(\mathbf{u}),D_{\mathbf{u}}(G)(\mathbf{u}),r)\geq N_{G(\mathbf{u}),\sigma}(0,r)-N^{(1)}_{G(\mathbf{u}),\sigma}(0,r).$
Write $G=\sum_{{\bf i}}a_{{\bf i}}{\bf x}^{{\bf i}}$ with $a_{{\bf i}}\in\mathbb{K}_{\mathbf{u}}$. Then, from the definition of $\mathbb{K}_{{\bf u}}$, there exists a nonnegative real $c_{G}$ (depending on $G$) such that
(4.8) $T_{G,\sigma}(r)\leq_{\rm exc}c_{G}\left(\mathfrak{X}^{+}_{\sigma}(r)+\log r+\max_{1\leq i\leq n}\{\log T_{u_{i}}(r)\}\right)$
after shrinking radii $r_{j}$ suffciently small (depending on $G$). Furthermore, by Lemma 13, we have
$T_{D_{\mathbf{u}}(G),\sigma}(r)$ $\leq 2T_{G,\sigma}(r)+\deg G\sum_{i=1}^{n}T^{M}_{u_{i}^{\prime}/u_{i},\sigma}(r)$
(4.9) $\leq_{\rm exc}(2c_{G}+n\deg G)(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i}}(r)\}).$
By Lemma 22 and (4.8), either $G$ and $D_{\mathbf{u}}(G)$ are coprime or or there exists a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})\in\mathbb{Z}^{n}$ with $\sum_{i=0}^{n}|m_{i}|\leq 2\deg G$ such that
$T^{M}_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\sigma}(r)\leq_{\rm exc}T_{G,\sigma}(r)\leq_{\rm exc}c_{G}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i}}(r)\}).$
It remains to consider where $G$ and $D_{\mathbf{u}}(G)$ are coprime. Let $\epsilon>0$ be given, which may assume to be sufficiently small. By apply Theorem 20, after shrinking $r_{j}$ small enough (depending on $\epsilon$, $\bf u$} and $G$), we find $m$ and $c_{i}>0$, $1\leq i\leq 4$, such that either
(4.10) $T^{M}_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\sigma}(r)\leq_{\rm exc}c_{1}(T_{G,\sigma}(r)+T_{D_{\mathbf{u}}(G),\sigma}(r))+c_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
@@
$N_{\rm gcd,\sigma}(G({\mathbf{u}}),D_{\mathbf{u}}(G)({\mathbf{u}}),r)$
(4.11) $\leq_{\rm exc}{\epsilon\over 2}\max_{1\leq i\leq n}T_{u_{i}}(r)+c_{3}(T_{G,\sigma}(r)+T_{D_{\mathbf{u}}(G),\sigma}(r))+c_{4}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
We can conclude the proof by combining (4.8), (4), (4.10), and (4). ∎